OpenStudy (anonymous):

Let me know if I have this one correct please: In 1991, the life expectancy of males in a certain country was 65.5 years. In 199, it was 69.1 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991. E(t)=_.7__t+_69.1_ (round to the nearest tenth). Use the function to predict the life expectancy of males in 2009. E(18)=_78.5__(round to the nearest thenth).

OpenStudy (anonymous):

you probably made mistake In 199, it was 69.1 years or it's really 199 year?

OpenStudy (anonymous):

oops, 1996 i believe..let me go check..

OpenStudy (anonymous):

yes, its 1996

OpenStudy (anonymous):

also what ___ means?

OpenStudy (anonymous):

thats just where i was insertinf my answers..ignore the ___

OpenStudy (anonymous):

i don't really understand. E(t)=_.7__t+_69.1_ (round to the nearest tenth). Use the function to predict the life expectancy of males in 2009. E(18)=_78.5__(round to the nearest thenth). this lines were given? or you wrote them?

OpenStudy (anonymous):

i had to insert my answers E(t)=___t+___ and E(18)=____, so my first anser to that problem is E(t)=.7t+69.1 and my second answer is E918)=78.5.

OpenStudy (anonymous):

i get E(t)=1.4t+1900 and second is 2009=1.4t+1900 t=77.9

OpenStudy (anonymous):

hmm, i don't know how i came up with my answer..1.4 is rounded to the nearest tenth??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[y=mx+b\] need to solve this system \[\begin{cases}1991=65.5m+b\\1996=69.1m+b\end{cases}\]

OpenStudy (anonymous):

just curious how you get the i get 1900, I am doing something wrong figuring my equation

OpenStudy (anonymous):

oh ok..

OpenStudy (anonymous):

well 1991 and 1996 are y1 and y2 while 65.5 and 69.1 are x1 and x2 so you have 2 points and need to find a function you can do it using various methods, idk if method posted by me is familiar to you

OpenStudy (anonymous):

no, but its looks easier than a text book..text book confuses

OpenStudy (anonymous):

I have one last problem and I am having a hard time with, can you help me again?

OpenStudy (anonymous):

ok, post a new question :)

OpenStudy (anonymous):

\begin{cases}0=65.5m+b\\5=69.1m+b\end{cases} you should solve this because it was said Let E represent the life expectancy in year t and let t represent the number of years since 1991.

OpenStudy (anonymous):

so if it's 1991 they mean t=0 or not?

OpenStudy (anonymous):

oh boy, now I'm confused...

OpenStudy (anonymous):

well, i don't understand that sentence since i am not native speaker, so you should help me to understand it :P

OpenStudy (anonymous):

let me see my notes again..

OpenStudy (anonymous):

this is what i have in my notes. 1991(65.5 years), 1996 (69.1 years). m=65.5-69.1/5years=3.6/5 3.6/5=0.72(round to nearest tenth) E(t)=.7t+65.5? The above is what I have for my notes! So here is where I get confused: E(18)= how do i figure to do this?

OpenStudy (anonymous):

yeah it was correct! sorry for confusing you ><

OpenStudy (phi):

Let's see if I can add to the confusion: year t exp 1991 0 65.5 1999 8 69.1 we have two points using t=x, and age=y: (0,65.5) and (8,69.1) we have 2 points, so find the equation of a line. One way is first find the slope. change in y / change in x = (69.1-65.5)/(8-0)= 3.6/8 =0.45 now use (y-y1)= m(x-x1) where (x1,y1) is one of your points. using (0,65.5) as (x1,y1) we get y-65.5= 0.45(x-0) and y= 0.45x+65.5

OpenStudy (anonymous):

now you need to put 18 instead of t

OpenStudy (anonymous):

phy it's inccorrect, it's 5-0 not 8-0

OpenStudy (anonymous):

hehe. here is what i got for 18: E(18)=0.72+65.5 12.96=65.5=78.46

OpenStudy (phi):

I took a guess at 199.

OpenStudy (anonymous):

you need to write E(t)=0.72t+65.5 E(18)=0.72*18+65.5=78.46

OpenStudy (anonymous):

because you put 18 instead of t

OpenStudy (anonymous):

sorry phi...its 1999

OpenStudy (anonymous):

and after approximation it's 78.5

OpenStudy (anonymous):

its 1999? or 1996? :P

OpenStudy (anonymous):

sorry again, 1996...

OpenStudy (anonymous):

so islandgirl you were correct at the beginning >< and i just fluttered everything

OpenStudy (anonymous):

sorry ok now lets move to your next question

OpenStudy (anonymous):

lol, no worries, we all learning here...ok on to the other problem.

OpenStudy (anonymous):

good try phi, great effort.. :)